OpenStudy (anonymous):

Multiplying Binomials: (x + 2) ( x - 9) EXPLAIN, please.

OpenStudy (anonymous):

Use FOIL. So F=first or x*x=x^2. O=outer or x*-9=-9x. I=inner or 2*x=2x. L=last or 2*-9. So we add everything together and get x^2-7x-18.